the dimensional formula of the permittivity of vacuum

Grade:12

2 Answers

Saurabh Kumar
7 years ago
PEMITTIVITY= (CHARGE)^2 /(FORCE)*(DISTANCE)^2 =[TA]^2 / [MLT^ -2] [L]^2 =[TA]^2 / [ML^3T^ -2] =[M^ -1 L^ -3 T^2+2 A^2] = [M^ -1 l^ -3 T^4 A^2]
Anjali kumari
5 years ago
It's dimension is M^-1L^-3T^4A^2. As permittivity=q1q2/4πFr^2. Sahi sell do aap devi sathi del zal zero aap so do LO